OpenStudy (anonymous):

The area of a circle is approximately 3.14 times the radius squared. Which of the following expressions is a correct way to write this, if the radius is 3x?

OpenStudy (anonymous):

3.14(3x)2 (3x)(3.14)2 3.14(3)x2 9(3.14)x answers

OpenStudy (anonymous):

The area is 3.14 times the radius squared, and the radius is 3x. So: \[A = \pi r^2 = \pi (3x)^2 = 9x^2 \pi \]